46 members of a band need to travel to a concert. No more than twelve members can ride in a van at one time. What is the minimum number to take the members of the band to the concert
Answer by Theo(13342)

46 / 12 = 3 plus a remainder of 10.
3 * 12 = 36
46 - 36 = 10
a minimum of 4 vans are required.
3 of them will carry 12 passengers.
1 of them will carry 10 passengers.